From GlueXWiki
Jump to: navigation, search

February 13, 2014 FDC meeting


  1. Installation [1],[2],[3],[4],[5],[6],[7],[8]
  2. Engineering (Bill)
  3. Electronics problematic channels (Chris, Nick)
  4. Mini-DAQ results FDC E-log (Beni)
  5. Other


Participants: Eugene, Fernando, Bill, Dave, Nick, Chris, Simon, Beni, and Lubomir.


- On Thursday last week we inserted the FDC into the magnet (see pictures attached). It was taken out of the magnet on Monday. The test was successful and we learned how to better handle the cables during insertion, what we discussed on the meeting. As Chris suggested, during the extraction, we used a strap to keep the cables up when moving the FDC. Bill suggested, for the final insertion, to use bearings (at least three) attached to the rails to be able to move the cables together with the rest, FDC and the mesh. Another discussion: how to attach the cables to the face of the magnet. On Monday we already exercised this at one quadrant using holding rings attached to the magnet. The main problem is that one has to spread the cables evenly around the periphery to make space for the CDC and also to be able to work on the BCAL modules. Beni proposed to extend the mesh further at the upstream side to cover completely the CDC, however the question was if there will be space to work around the BCAL modules. Bill proposed to build a ring (attached to the platform) with the same diameter as the mesh to hold the cables and, at the same time, be able to work on the other detectors. Fernando proposed to build a bigger movable ring attached to the face of the magnet, to support the cables.

- Before the insertion, the cooling system failed again, the reasons are unknown so far, but there was no LV at this time. During the insertion the Fluorinert houses were disconnected and taped with duck tape at the chiller side (due to the lack of proper caps), as a result we lost >1 gal of Fluorinert: need proper caps for the final insertion!

- Tom informed us that the power in the Hall and the CC will be OFF for 3-4 days starting February 20 (next Thursday). For the FDC we need the gas running all the time and will have to replace the mass flow controller (which requires power) with a rotameter. To avoid potential high pressure in the chamber with such a system Beni suggested installing a bubbler at the gas input.


- Bill will travel to Hexagon Metrology in Charlotte, NC next Wednesday/Thursday to scan the cathode.

- We need to modify the mesh to make sure there are no closed loops and also to insulate it from the FDC packages (now it's connected through two metal plates). The mesh will have a separate ground. Bill will work on that.

- Dave started working on the hardware shut off of the LV crate from the chiller. He needs to identify how exactly to interpret the signals about the chiller status and then will figure out with Fernando what the PCB should be doing. The board can be mounted on the same rack (back side) where the LV crate is. As for the MPOD controller settings Dave will talk to Hovanes.


- Fernando did noise/signal measurements on the FDC: good news, the noise at/around FDC is low: 8mV peak-to-peak. Will report at the collaboration meeting. Eugene: expect many more new devices to be turned on in the Hall.

- Lubomir prepared a table (linked above) with the problematic channels found so far with a scope or, for the wires, with the Beni's mini-DAQ. The problems are grouped in several types:

  1. conductive tape issues (or related to that): 13 channels found so far with a scope and the shaper board (doesn't see the inner 8 channels), some were fixed already, half of the problems are next to the cooling manifolds where it's difficult to bend the tubes.
  2. two groups of four pre-amps connected to the same LV cable showing similar problem: one of the channels saturates and the signals on all the channels of the other cards have a long (500ns) tail. Fernando will look at it, but suggested first replacing the cable, then the pre-amps.
  3. ~10 cards with cable issues, meaning if you replace the cable it works. Nick and Chris will come tomorrow and fix the bad connectors at the detector side.
  4. 6 wire cards with mostly one channel not working as found by the Beni's tests. Most of these cards are at the same position at the four packages (C1,S1-24) and these channels were working before the FDC insertion. Chris: we had to use force when inserting the FDC and may have pulled on the whole bundle that connects the cables at the same positions at the packages. All these channels have been fixed by pushing the cards back in the pre-amp connector.
  5. 4 cards all in the first package and always at the same place w.r.t to the cell: S49-72 with 5 to 10 first channels showing 1-30kHz noise at 8V threshold. The noise goes down to 10-500Hz at 10V. The most interesting is that the noise is there even if turn off the HV and that there's no noise when looking with an analog pre-amp. Fernando will look into this.
  6. 2 cards with very small (in amplitude) LVDS signals. Fernando: check cables, replace pre-amps.

- All the negative (field) HV channels have been checked by Anatoly measuring the capacitance at the electronics end of the HV cable. It ranges from 1.9 to 2.5 uF. If the cable is disconnected from the detector side the capacitance is 1.2 uF. We conclude the negative HV connections are OK.


- Beni presented his latest results from the mini-DAQ (E-log entries 138-141). Beni has now all the crates (four) with TDCs read out so we can test all the wires at the same time. The plots show the problems discussed above.

- Beni proposed to swap the top and bottom cables as connected to each of the TDCs, so that "it makes sense", and he did it. Now there's correspondence between the detector and the TDC channel numbering. If everybody agrees we will have to re-label the connectors and change the translation table.